OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (0, 0) and (3, 30). 10 0 5 −10

OpenStudy (anonymous):

slope= (y2-y1)/(x2-x1) slope=(30-0)/(3-0) slope=30/3 slope= 10